• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name vasulk

    Log entry time 08:07:41 on February11,2003

    Entry number 96045

    keyword=Owl Shift Summary

    Begin shift by moving BHWP IN, Kin 3.7.3 at 0 degrees
    RWP 6250, PC-pos: 7.704, PC-neg: 7.538

    P0 = 2.15 GeV for both Right and Left Arm
    00:00-04:40 Helicity in delay =0 mode, Pair, toggle

    00:07 Start Kin 3.7.3, BHWP IN
    01:40 Finish Kin 3.7.3, BHWP IN

    01:43 Rotate holding field to 270 degrees
    01:49 At 270 degrees with Lasers on

    01:51 Begin Kin 3.7.5, BHWP IN, helicity delay = 0
    03:25 Beam down, MCC will call back and give an update
    03:33 MCC said there is a bad RF zone, ~ 1 hour to fix
    04:05 Update, Experts called in to look into RF problem ~ 40 minutes more

    Decide to end Kin 3.7.5, BHWP IN, helicity delay =0 early

    04:06 Change momentum on both Arms to 1.987 GeV/c
    04:10 Begin field rotation to 0 degrees
    04:13 Left Arm momentum set, Right dipole is still settling
    04:17 At 0 degrees, target moved to BeO position
    04:37 Right Arm momentum set
    04:40 MCC calls and are getting ready to restore beam

    Asked for them to restore G0 helicity settings
    Delay = 8 cycles, quartet pattern, random mode

    04:50 Beam back, spot and OTR look good
    04:55 RF zone down again, ~ 15 mintues to fix this time
    05:15 MCC ready to restore beam
    05:37 Start N2 reference cell run, pressure 149 psig
    05:53 End N2 reference cell run, pressure 149 psig
    Finish 3 GeV program, ask for energy change to 5 GeV

    06:00 Change Left and Right arm momentum to 2.379 GeV/c
    Cycle Q2 and Q3 from both spectrometers
    06:05 R-Arm Q2 tripped

    06:11 NMR measurement: 37.2%

    06:15 L-Arm momentum set, R-Arm dipole still settling, and
    R-Arm Q2 cannot be reset remotely
    06:20 Talked to Mark Stevens, I will make an access to reset Q2
    06:35 R-Arm dipole set, I will need to wait while MCC changes
    energies before I can make an access

    06:47 L-Arm moved to 32 degrees
    07:00 MCC called, they are sending tuned beam to check orbits
    07:05 R-Arm moved to 32 degrees
    07:15 Control Acces to reset Q2 and target computer
    07:40 Control Access complete, Q2 cycled to 1400.0 Amps
    07:42 R-Arm momentum ready, waiting for beam

    Run Summary

    Kin 3.7.3, BHWP IN, 0 degrees, helicity delay = 0

    run # 20931/1931: from 00:07 to 01:06
    run # 20932/1932: from 01:09 to 01:40

    Kin 3.7.5, BHWP IN, 270 degrees, helicity delay = 0

    run # 20933/1933: from 01:51 to 02:55
    run # 20934/1934: from 02:58 to 03:34 (last 9 minutes no beam)

    Kin 3.6

    run # 20935/1935: beam spot and raster check
    run # 20936/1936: empty reference cell run
    run # 20937/1937: N2 reference cell run

    Cosmic run # 20398/1938